OpenStudy (anonymous):

What is the greatest common factor of 35x6 and 7x2? Select one of the options below as your answer: A. x^2 B. x^6 C. 5x^2 D. 7x^2

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

35*6=5*7*2*3 7*2 G.C.D.=Common factors in both=7*2